HUBBLE SPACE TELESCOPE DAILY REPORT #2579 PERIOD COVERED: 0000Z (UTC) 03/13/00 - 0000Z (UTC) 03/14/00 Daily Status Report as of 074/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally. 1.2 Completed Two Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.3 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 1.4 Completed STIS/CCD/MA1 8613 (Coordinated Observations of Stellar Flares on AD Leo) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to obtain high resolution spectra of the dMe star AD Leo during periods of quiescence and during flares to study the physics of atmospheric heating in flare stars. The proposal completed with no reported problems. 1.5 Completed Three Sets of WF/PC-2 8442 (WF/PC-2 Cycle 8 Standard Darks) The WF/PC-2 was used to make a series of Cycle 8 standard dark calibration observations. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of WF/PC-2 8457 (UV Earth Flats) The WF/PC-2 was used to obtain a series of Ultraviolet Earth flats. The observations were completed with no reported problems. 1.7 Completed WF/PC-2 8436 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to make archive observations parallel to the prime science observations as part of a POMS test proposal. This is a POMS test proposal designed to simulate scientific plans. The observations were completed as planned, and no anomalies were reported. 1.8 Completed STIS/CCD 8327 (Observations of Eta Carinae: The Central Star The Space Telescope Imaging Spectrograph (CCD) was used to observe Eta Carinae {1} to obtain the spectrum of the star at ``normal'' times; {2} to extend the wavelength coverage below 1700 Angstrom; {3} to seek evidence for a hypothetical companion star; {4} to learn whether yhe star varies significantly when it is not near a discrete spectroscopic event; and {5} to study the bizarre spectra of dense, slow-moving, extremely peculiar ejecta located 0.15" to 1.5" from the star. The proposal completed nominally. 1.9 Completed WF/PC-2 8239 (Imaging the NLR in a Complete Sample of z<0.5 Radio-Quiet PG Quasars) The WF/PC-2 was to be used to observe quasar PG1012+008 and to determine its role in the evolution of galaxies, where a central engine leaves its visible imprints in the galaxy through strong ionization and heating of the ISM. As detailed in 2.1 and HSTAR 7535, the acquisition and re-acquisitions for this proposal failed, affecting all 14 observations. 1.10 Completed STIS/CCD 8796 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation IV) The Space Telescope Imaging Spectrograph (CCD) was to be used in the parallel mode to make some low galactic latitude archive observations. As detailed in 2.1 and HSTAR 7535, the acquisition for this proposal failed, affecting one observation. 1.11 Completed WF/PC-2 8338 (The Evolution of the Host Galaxies of Powerful Radio Sources) The WF/PC-2 was used to make observations of the host galaxy of the powerful radio source known as MRC2028-29. A combined data set of radio galaxies and quasars will allow powerful tests of the relationship between radio galaxies and quasars and the evolution of the host galaxies of powerful radio sources over a wide range of look-back times. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 8 Successful: 7 Scheduled Re-acquisitions: 6 Successful: 4 Per HSTAR 7535, the acquisition at 073/145325Z failed when the search radius limit was exceeded for FGS-3. The succeeding re-acquisitions at 073/162841Z and 073/181059Z also failed. He proposals described in 1.9 and 1.10 were affected. 2.2 FHST Updates: Scheduled: 20 Successful: 19 The full maneuver updates at 073/185415Z and 073/185700Z failed on FHST-1. HSTAR 7536 was written. The subsequent acquisition was successful, however. HSTAR 7537 documents that the roll delay update scheduled for 073/232248Z failed due to tracker 1. 2.3 Operations Notes: Other than the acquisition and update problems detailed above, there were no reported operational problems. Returned Battery 3 to the normal six-battery configuration at 073/2230Z, following the completion of the Battery 3 Capacity Check Test. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW